HUBBLE SPACE TELESCOPE DAILY REPORT #2132 PERIOD COVERED: 0000Z (UTC) 05/27/98 - 0000Z (UTC) 05/28/98 Daily Status Report as of 148/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed NIC/1/2/3 7875 (Searching for Old Stellar Populations in Candidate Proto-Spheroidals) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations searching for old stellar populations in candidate proto-spheroid SA57-7042. We have compelling evidence from WF/PC-2 images that these galaxies have optical colors gradients. The near-infrared is necessary, however, to reveal color gradients large enough to separate young and old stellar populations. Only NICMOS imaging provides enough spatial resolution and depth to resolve these gradients. The observations were completed as planned, and no problems were reported. 1.4 Completed Five Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.5 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.6 Completed NIC/1/2/3 7372 (Fundamental Plane, Morphology-Density Relation, and Lensing in the z=0.83 Cluster) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of galactic cluster Gal-Clus-1. These data will be used to augment the HST mosaics of the clusters with extensive ground-based spectroscopy and imaging at the Keck, MMT, NTT, and WHT telescopes. The observations were completed as planned, and no problems were reported. 1.7 Completed WF/PC-2 7590 (Cosmological Parameters Omega and Lambda From High-Redshift Type IA Supernovae) The WF/PC-2 was used to make observations of type 1A supernova 14H0.9Z2 in order to study its redshift. The HST was used to make direct measurements of the cosmological parameters Omega, Lambda, and curvature using Type 1A supernovae as a calibrated "standard candle". The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Two Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.9 Completed WF/PC-2 7336 (Cosmological Parameters Omega and Lambda From High-Redshift Type Ia Supernovae) The WF/PC-2 was used to make observations of the high redshifted Type Ia supernovae SN94-8H-Z0, SN91-5H-Z0 and 11H0.5Z1. HST can make a direct measurement of the cosmological parameters, Omega, Lambda, and curvature, using Type Ia supernovae as calibrated "standard candles." The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed STIS/CCD/MA1/MA2 7673 (STIS MAMA Sensitivity and Focus Monitor) The Space Telescope Imaging Spectrograph (CCD/MA1/MA2) was to be used to make observations of standard star GRW+70D582 in order to perform a MAMA sensitivity and focus monitoring program. This is to monitor sensitivity of each MAMA grating mode to detect any change due to contamination or other causes. As documented in HSTAR 6494 and described in 2.1, the guide star acquisition for this proposal failed and the take data flag remained down. 1.11 Completed NIC/1 7867 (NICMOS Imaging of the Dusty CfA Seyfert Nuclei) The Near Infrared Camera (Camera One) was used to make observations of the Seyfert Galaxy NGC-5273. These data will allow us to determine if the Seyfert Type 1/2 dichotomy is best explained by the unified model, the evolutionary picture, or a combination of both of these scenarios. The observations were executed as scheduled, and no anomalies were reported. 1.12 Completed NIC/2 7887 (A Survey of Gravitational Lenses as Cosmological Tools II) The Near Infrared Camera (Camera Two) was used to make observations of the gravitational lens candidate system UM425. The propose is to extend our IR survey of gravitational lenses and gravitational lens candidates by obtaining NICMOS images of 14 additional lens systems or candidates. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 7 The acquisition at 147/183847Z ultimately failed to RGA control due to failed V1 FGS attitude error vector checks. The acquisition initially succeeded, but a subsequent loss of lock occurred. Recovery from that event was unsuccessful. HSTAR 6494 was written. The STIS proposal described in 1.10 was affected. Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 21 Successful : 21 2.3 Operations Notes: As a result of the acquisition failure described in 2.1, HSTAR 6495 was written, documenting three STIS status buffer error messages. The flags were cleared as directed by a ROP. The EDAC error counts were cleared at 147/1855Z, per ROP SR-1. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH